# SF-R You Sure? The Conflicting Role of Star Formation Rates in Constraining the Evolution of Milky Way Analogues in Cosmological Simulations

### v1

@@ 5.1 Observationally selected analogues
These weights, , now define our OSAs at redshift . Note our choice here to use the original weights () and not the adjusted weights () since we are dealing with the MW-ness of individual galaxies as they evolve through time (the criterion we defined for application of the original method), rather than the MW-ness of a population (the criterion we defined for application of the adjusted method). This process is repeated at each snapshot in the simulation.
@@ 5.2 Tracing observationally selected analogues forwards in time
It is worth noting again that the “pseudo selection box” in Figure 5 increases in size with each previous snapshot222We stress the adjective “pseudo” as the OSAs in this study are indeed selected using the weights discussed in Section 3.2.2 rather than the traditional box cut method, and we say “selection box” here only as an illustrative tool.. This indicates that with increasing redshift, more and more galaxies are being selected as OSAs that are not part of the MWAs(0) group – in other words, many OSAs do not become MWAs today. We introduce the term contamination to describe this concept of non-MWAs(0) galaxies (MWA-“imposters”) being considered OSAs at higher redshift.
@@ 6.1 Returning to the guiding questions
What defines a Milky Way analogue? Since this work focused on studying the SFHs of MWAs, we have chosen to select our MWAs on stellar mass (SM, ) and star formation rate (SFR, ) based on the values found for the MW by LNB15: and . Our selection was determined using a distribution around the MW’s SM and SFR (Equation (4)), which was then re-weighted against the background sample by normalization via a pdf to account for the overwhelming skew towards low-SM or high-SFR galaxies. This method of MWA selection resulted in a set of weights, called , that provide a metric of similarity between each galaxy in the simulations to the MW itself. In addition to our MWAs, we have identified two control groups, one selected on SM alone and one selected on SFR alone, in order to answer question (2).
@@
It seems that in terms of SFR, selecting only on SM is more likely to pick out MWAs at early times, but selecting only on SFR is more likely to pick out MWAs at late times. Further, this crossover from “early” to “late” occurs much earlier in TNG than in EAGLE, and the difference between the MWAs and the SFR-controlled sample is also much larger in TNG than in EAGLE. In fact, the SFHs of the EAGLE MWAs only diverge from the SM-controlled group at a point in their evolution at which they have already formed the vast majority of their stars, and so the elevated SFR at late times has a negligible effect overall. Considering EAGLE’s success in reproducing the MW’s red colour and TNG’s too-blue MWAs in Figure 4 caused by lack of quenching, we may be more inclined to trust EAGLE’s SFHs. In this case, present-day SM alone may be “good enough” to identify MWAs at higher redshift.
@@ 6.2 Sliding the “selection box” with shifted analogues
To further investigate the restrictive effect of a small SFR window, we move our SFR “pseudo selection box” at to higher and lower SFR values and study the evolutionary histories of the resulting new groups of galaxies. In other words, rather than studying MWAs – a group of galaxies that converge to the MW’s SM and SFR today – we study a group of galaxies that converge to the MW’s SM today, but to a different value for present-day SFR. We consider two groups that end with higher SFR than the MW and two groups that end with lower SFR in both simulations. We keep the size of our selection window the same for these new groups (i.e., we keep the same uncertainty in SFR as we used to identify our original MWAs), and again use the adjusted method (Section 3.2.3).
@@
shows the SFHs and SMHs for these new galaxy groups – the shifted analogues – in violet, from lowest present-day SFR selection (which we call SFR(0)) on the left to highest on the right. The original MWAs for both simulations are shown in the centre panels as green for EAGLE and orange for TNG. The solid lines, shaded regions, and black and blue colouring are defined the same as in Figure 6; however, the SFR-controlled samples have been adjusted to converge to the same SFR(0) as the shifted analogue groups, while the SM-controlled samples remain the same for all panels. The exact values of SFR(0) used for each shifted analogue group as well as the effective sample sizes (calculated with Equation (6)) for all groups are shown in the respective figure panels. We see that with increasing SFR(0), the MWAs evolve less closely with the SM-controlled group and more closely with the SFR-controlled group.
@@
As a general trend in the SMHs for both simulations, the higher the SFR(0), the higher the overall and final SM of the SFR-group, but the lower the overall and final SM of the MWAs. In Figure 10, this is seen as increasing SFR(0) corresponds to raising of the SFR-controlled group (blue) and lowering of the shifted analogue group (violet). As SFR(0) increases, the shifted analogue group pulls further away from the SM-controlled group above and falls closer towards the SFR-controlled group below (with the exception of the leftmost panel for EAGLE), indicating that SFR becomes a more important selection criteria for these shifted analogues. In fact, on the highest SFR(0) end, the shifted analogue SMHs line up more closely throughout time with the SFR-controlled group, rather than tracing the path of the SM-controlled group as they do on the lowest SFR(0) end (second lowest SFR(0) for EAGLE).
@@
The result of this trend tied to higher SFR(0) is that in both simulations, to produce galaxies with SM fixed to be similar to the MW at present day, as in the shifted analogue group, galaxies with higher present-day SFR (rightmost panels) must assemble their mass more slowly in order to counteract this late-time elevated SFR as opposed to galaxies with lower present-day SFR (leftmost panels). However, when present-day SM is not fixed, as in the SFR-controlled group, galaxies with higher SFR(0) exhibit higher present-day SM and assemble their mass earlier. This is to be expected, as galaxies that have been high star forming throughout their evolution should have more mass in stars.
@@
Considering now the SFHs, we see that extreme values on either the high or low end for SFR(0) result in a more significant divergence from the SM-controlled group in the shifted analogue evolution. The highest SFR(0) selection in both simulations shows the shifted analogue groups diverging from the SM-controlled groups very early and following more closely the path of the elevated SFR-controlled group. In the case of the lowest SFR(0), the two simulations produce differing SFHs – in both simulations, the SFR of the shifted analogues starts high, peaks early, and then rapidly quenches; the difference comes in the comparison between the shifted analogues and the SM-controlled sample. In EAGLE, the depressed SFR-controlled group pulls the shifted analogues below the SM-controlled group; while in TNG, the shifted analogues align nearly perfectly with the SM-controlled group, further signifying the intensity of quenching in this particular simulation.
@@
Taking a step back from MWAs, we consider for a moment how SM and SFR influence a galaxy’s evolution more generally. Collectively, the SMHs of the varying SFR-controlled samples (i.e., galaxies without a fixed final SM) demonstrate that galaxies with high present-day SFR most likely also have high SM (if we consider the MW to have high SM) and galaxies with low present-day SFR most likely also have low SM. Now looking at the SFHs, in both cases of high or low present-day SFR, these galaxies are currently and have always been forming stars. However, turning to the shifted analogues (i.e., now fixing present-day SM to be high), only galaxies also fixed to have high SFR(0) will have elevated SFHs, and galaxies fixed to have lower SFR(0) will be forced to peak early – as they must be high star forming at some point in order to end up with high mass in stars – and then undergo rapid quenching. Thus, it appears that a fixed present-day SM correlates to whether a galaxy exhibits quenching, as galaxies seem not to want to quench without this constraint, and a fixed present-day SFR correlates to a galaxy’s final mass bracket.
@@
Finally, we note that when comparing simulations, varying SFR(0) results in similar evolutionary tracks in the shifted analogues, but different conclusions regarding the significance of SFR as a selection parameter compared to SM alone. In EAGLE, the extreme choices in SFR(0) show drastic differences in evolutionary tracks of the shifted analogues – on the low SFR(0) end, the SM assembly is too quick and star formation quenches too rapidly compared to the SM-controlled group, and on the high SFR(0) end the SM assembles too slowly and the SFR does not quench enough compared to the SM-controlled sample. However, the original MWAs (centre panel, green) trace closely the SM-controlled group for both the SMHs and SFHs as previously discussed, and lay in a sweet spot where present-day SFR and SM seem to correlate well, implying SFR may be a redundant selection criterion.
@@
In contrast, the extreme low end of SFR(0) (leftmost panel) in TNG results in near perfect overlap between the shifted analogues and the SM-controlled group, while the original MWAs (centre panel, orange) assemble their SM too slowly and avoid strong quenching as compared to the SM-controlled group. The MWAs do not lay in the same sweet spot as in EAGLE, and present-day SFR is necessary to include as a parameter in MWA selection.
@@ 6.3 Re-evaluating the selection parameters
The results of our study of MWA SFHs in EAGLE and TNG produced conflicting views on the importance of present-day SFR as a MWA selection criterion. The simulations imply different conclusions in the context of MWA evolutionary tracks – in EAGLE, SM is sufficient a stand-alone selection criterion for MWAs, but in TNG it is necessary to include SFR as a parameter. Both simulations do agree, however, that including SFR as a parameter for OSAs drives down selection efficiency.
@@
If we are to better trust EAGLE’s SFHs – since they more accurately reproduce the MW’s colour – that SFR can safely be neglected as a selection parameter, then this combined with our result that the SFR constraint in observational studies may not be useful suggests that the inclusion of SFR as a selection parameter may be too restrictive.
@@
To test the effect of loosening the SFR constraint, we have re-calculated , as in Equation (10), using a series of expanding SFR “pseudo selection boxes”, corresponding to a change in SFR uncertainty ranging from 1 to 3 times the value of 0.19 provided by LNB15 in regular intervals of 0.2. Here we remind the reader that the use of the term “pseudo” is to signify that neither the MWAs nor the OSAs are identified using a harsh selection box, but are weighted with respect to this outlined uncertainty. The results are plotted in Figure 11,
@@
where increasing transparency in the line corresponds to increasing size in the SFR “pseudo selection box”. We discover that loosening the SFR constraint in this way significantly reduces contamination at higher redshifts. Increasing the SFR cutoff produces a smoother transition in going back in time, particularly for the largest selection box which produces selection nearly as efficient, and sometimes even better, than the SM-controlled group. Future studies of MWAs, especially in observations, may benefit from such an expanded selection window in SFR. We also note that an expanded selection window should have minimal consequences on our study of MWA evolution, as the effects of incorporating slightly higher and lower SFRs should essentially cancel, as evident from the panels directly to the right and left of the MWAs in Figure 10.
@@
To understand why expanding the “pseudo selection box” in SFR increases selection efficiency, we first explain why for the SFR-controlled group in Figure 9 is so low, which ends up being simulation-dependent. In EAGLE, instantaneous SFR tends to bounce around between snapshots due to high SFR variability in the simulation (see Figure 7), pushing the limits of the MWA distribution further out and expanding the cutoff of the OSA “pseudo selection box”. Consequently, many non-MWA(0) galaxies will get picked up as OSAs, driving down.
@@
There is a different cause for the decline in in TNG. We previously noted that for the MWAs in Figure 7 plateaus after going back Gyr, indicating that SFR variability is not as strong as in EAGLE; rather, the MWAs as a group evolve cohesively. This is illustrated in Figure 10, where all groups but the one selected on the lowest SFR(0) exhibit little dispersion in their SFHs. What is causing the drop in TNG then is the overlap between these shifted analogue groups. Looking at the SFHs of the shifted analogues to either side of the MWAs, we see that these groups all occupy much of the same space. Thus, the OSA “pseudo selection box” is picking up many of these galaxies with SFR(0) slightly above or below the MW’s present-day value, hence why expanding the SFR cutoff significantly increases efficiency.
@@ 7 Summary & conclusions
Our galaxy population comes from the cosmological, hydrodynamical simulations EAGLE and TNG. We identify MWAs from this greater population using a distribution centred on the MW’s SM and SFR given by Equation (4), re-weighted against the background galaxies to offset a skew towards low SM or high SFR. This process results in a set of MW-ness weights, , given by Equation (7), providing a metric of similarity to the MW for each galaxy within the simulations. In addition to our MWAs, we use the same method to identify two control groups, one selected on SM alone and the other on SFR alone, in order to determine the role of each parameter in altering the sample of MWA evolutionary tracks.
@@
The main results of our work are summarized as follows:
@@
MWAs’ mass assembly histories are best predicted by present-day SM. The SMHs of the MWAs in Figure 6 (right panels) follow very closely to the SM-controlled group, with near-perfect overlap between the two groups in EAGLE, indicating all galaxies with a present-day SM comparable to the MW assemble their mass in stars similarly over time, regardless of final SFR. From this analysis we also learn that MWAs generally assemble a lot of SM relatively quickly.
MWAs show a characteristic SFH profile predicted by present-day SM at early times and present-day SFR at late times, corresponding to an early peak in SFR that remains elevated through cosmic time. In Figure 6 (left panels), the SFHs of the MWAs align more closely with the SM-controlled group at early times, before crossing over to aligning more closely with the SFR-controlled group at late times. This crossover point, where the tracer of star formation evolution changes from SM to SFR, occurs much earlier and results in a stronger divergence in TNG than in EAGLE.
The divergence between the MWAs and SM-controlled sample mentioned above leads to another result: in TNG, MWAs do not quench as strongly as most galaxies of the same present-day SM. The SFHs in Figure 6 (bottom left panel) show that galaxies with a MW-like SM (i.e., the SM-controlled group) in TNG rapidly quench early in their evolution, but the MWAs are the ones “lucky” enough to bypass this shutoff of star formation. The reason for this stark difference when including SFR in the selection criteria for TNG, allowing them to avoid quenching, will be the focus of a future publication. We also note that this lack of quenching is why the TNG MWAs appear too blue in Figure 4, as the galaxies are forming new stars at and hence do not turn red.
In contrast to the above point, all MW-SM galaxies in EAGLE experience slow quenching with the MWAs holding only a slightly elevated present-day SFR in comparison (see Figure 6, bottom right panel). Thus, it seems again that present-day SM is the best predictor for MWA evolution in EAGLE. In fact, we conclude that present-day SM may be sufficient as a stand-alone selection parameter for MWAs in EAGLE, however in TNG it is necessary to include present-day SFR. We note that based on EAGLE’s success in reproducing the MW’s colour in Figure 4, we may be more inclined to believe EAGLE’s subgrid prescriptions and suggest SFR may be a redundant selection parameter.
@@
The authors would like to thank Sina Babaei Zadeh, Ishika Bangari, Eshal Arshad, Matthew Kustec, Cissy Kuang, and Zehao Peng for helpful conversations and research that provided new insights and improved the quality of this work, as well as Vivian Yun Yan Tan for comments on the manuscript. AMS would like to thank Aryanna Schiebelbein-Zwack, Anika Slizewski, and Braden Gail for their support and valuable comments, as well as Bart Ripperda for guidance and advice in the preparation of this work. JSS would like to thank Rebecca Bleich for her support for the continued existence of the Milky Way.
AMS was supported by funding from the Dunlap Institute, a Queen Elizabeth II/Walter John Helm Graduate Scholarship in Science and Technology from the Government of Ontario, Walter C. Sumner Memorial Fellowship from the University of Toronto, and a Canada Graduate Scholarships – Master’s (CGS-M) Award and Postgraduate Scholarships - Doctoral (PGS-D) Award from the Natural Sciences and Engineering Research Council of Canada (NSERC). JSS was supported by funding from the Dunlap Institute, an NSERC Banting Postdoctoral Fellowship, NSERC Discovery Grant RGPIN-2023-04849, and a University of Toronto Connaught New Researcher Award. JTM was supported by funding from an NSERC Banting Postdoctoral Fellowship, the Dunlap Institute, and the Canadian Institute for Theoretical Astrophysics (CITA). NWM acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC RGPIN-2023-04901). This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. KGI was supported by funding from the Dunlap Institute and from NASA through a Hubble Fellowship.

### v2

@@ 5.1 Observationally selected analogues
These weights, , now define our OSAs at redshift . Note our choice here to use the original distance weights () and not the population weights () since we are dealing with the MW-ness of individual galaxies as they evolve through time (the criterion we defined for application of the original distance method), rather than the MW-ness of a population (the criterion we defined for application of the population method). This process is repeated at each snapshot in the simulation.
@@ 5.2 Tracing observationally selected analogues forwards in time
It is worth noting again that the “pseudo selection box” in Figure 5 increases in size with each previous snapshot222We stress the adjective “pseudo” as the OSAs in this study are indeed selected using the distance weights discussed in Section 3.2.2 rather than the traditional box cut method, and we say “selection box” here only as an illustrative tool.. This indicates that with increasing redshift, more and more galaxies are being selected as OSAs that are not part of the MWAs(0) group – in other words, many OSAs do not become MWAs today. We introduce the term contamination to describe this concept of non-MWAs(0) galaxies (MWA-“imposters”) being considered OSAs at higher redshift.
@@ 6.1 Returning to the guiding questions
What defines a Milky Way analogue? Since this work focused on studying the SFHs of MWAs, we have chosen to select our MWAs on stellar mass (SM, ) and star formation rate (SFR, ) based on the values found for the MW by LNB15: and . Our selection was determined using a distribution around the MW’s SM and SFR (Equation (4)), which was then re-weighted against the background sample by normalization via a pdf to account for the overwhelming skew towards low-SM or low-SFR galaxies. This method of MWA selection resulted in a set of weights, called , that provide a metric of similarity between each galaxy in the simulations to the MW itself. In addition to our MWAs, we have identified two control groups, one selected on SM alone and one selected on SFR alone, in order to answer question (2).
@@
It seems that in terms of SFHs, selecting only on SM is more likely to pick out MWAs at early times, but selecting only on SFR is more likely to pick out MWAs at late times. Further, this crossover from “early” to “late” occurs much earlier in TNG than in EAGLE, and the final difference between the MWAs and the SFR-controlled sample is also much larger in TNG than in EAGLE. In fact, the SFHs of the EAGLE MWAs only diverge from the SM-controlled group at a point in their evolution at which they have already formed the vast majority of their stars, and so the elevated SFR at late times has a negligible effect overall. In this case, present-day SM alone may be “good enough” to identify MWAs at higher redshift.
@@ 6.2 Sliding the “selection box” with shifted analogues
To further investigate the restrictive effect of a small SFR window, we move our SFR “pseudo selection box” at to higher and lower SFR values and study the evolutionary histories of the resulting new groups of galaxies. In other words, rather than studying MWAs – a group of galaxies that converge to the MW’s SM and SFR today – we study a group of galaxies that converge to the MW’s SM today, but to a different value for present-day SFR. We consider two groups that end with higher SFR than the MW and two groups that end with lower SFR in both simulations. We keep the size of our selection window the same for these new groups (i.e., we keep the same uncertainty in SFR as we used to identify our original MWAs), and again use the population weights method (Section 3.2.3).
@@
shows the SFHs and SMHs for these new galaxy groups – the shifted analogues – in pink, from lowest present-day SFR selection (which we call SFR(0)) on the left to highest on the right. The original MWAs for both simulations are shown in the centre panels as green for EAGLE and orange for TNG. The solid lines, shaded regions, and black and blue colouring are defined the same as in Figure 6; however, the SFR-controlled samples have been adjusted to converge to the same SFR(0) as the shifted analogue groups, while the SM-controlled samples remain the same for all panels. The exact values of SFR(0) used for each shifted analogue group as well as the effective sample sizes (calculated with Equation (6)) for all groups are shown in the respective figure panels. We see that with increasing SFR(0), the MWAs evolve less closely with the SM-controlled group and more closely with the SFR-controlled group.
@@
As a general trend in the SMHs for both simulations (panels (f)–(j) for EAGLE, panels (p)–(t) for TNG), the higher the SFR(0), the higher the overall and final SM of the SFR-group, but the lower the overall and final SM of the shifted analogues. In Figure 10, this is seen as increasing SFR(0) corresponds to raising of the SFR-controlled group (blue) and lowering of the shifted analogue group (pink). As SFR(0) increases (from left to right), the shifted analogue group (pink) pulls further away from the SM-controlled group (black) above and falls closer towards the SFR-controlled group (blue) below, indicating that SFR becomes a more important selection criteria for these shifted analogues. In fact, on the highest SFR(0) end (the rightmost panels (j) and (t)), the shifted analogue SMHs line up more closely throughout time with the SFR-controlled group, rather than tracing the path of the SM-controlled group as they do on the lower SFR(0) end (lowest SFR(0) for TNG in panel (p), second lowest SFR(0) for EAGLE in panel (g)).
@@
The result of this trend tied to higher SFR(0) is that in both simulations, to produce galaxies with SM fixed to be similar to the MW at present day, as in the shifted analogue group (pink), galaxies with higher present-day SFR (rightmost panels (j) and (t)) must assemble their mass more slowly in order to counteract this late-time elevated SFR as opposed to galaxies with lower present-day SFR (leftmost panels (f) and (p)). However, when present-day SM is not fixed, as in the SFR-controlled group (blue), galaxies with higher SFR(0) exhibit higher present-day SM and assemble their mass earlier. This is to be expected, as galaxies that have been high star forming throughout their evolution should have more mass in stars.
@@
Considering now the SFHs (panels (a)–(e) for EAGLE, panels (k)–(o) for TNG), we see that extreme values on either the high or low end for SFR(0) result in a more significant divergence from the SM-controlled group in the shifted analogue evolution. The highest SFR(0) selection in both simulations (panels (e) and (o)) shows the shifted analogue groups (pink) diverging from the SM-controlled groups (black) very early and following more closely the path of the elevated SFR-controlled group (blue). In the case of the lowest SFR(0) (panels(a) and (k)), the two simulations produce differing SFHs – in both simulations, the SFR of the shifted analogues starts high, peaks early, and then rapidly quenches; the difference comes in the comparison between the shifted analogues and the SM-controlled sample. In EAGLE (panel (a)), the depressed SFR-controlled group pulls the shifted analogues below the SM-controlled group; while in TNG (panel (k)), the shifted analogues align nearly perfectly with the SM-controlled group, further signifying the intensity of quenching in this particular simulation.
@@
Taking a step back from MWAs, we consider for a moment how SM and SFR influence a galaxy’s evolution more generally. Collectively, the SMHs of the varying SFR-controlled samples (i.e., galaxies without a fixed final SM; blue) demonstrate that galaxies with high present-day SFR most likely also have high SM (if we consider the MW to have high SM; panels (j) and (t)) and galaxies with low present-day SFR most likely also have low SM (panels (f) and (p)). Now looking at the SFHs, in both cases of high (panels (d), (e), (n), (o)) or low (panels (a), (b), (k), (l)) present-day SFR, these galaxies are currently and have always been forming stars. However, turning to the shifted analogues (i.e., now fixing present-day SM to be high; pink), only galaxies also fixed to have high SFR(0) will have elevated SFHs, and galaxies fixed to have lower SFR(0) will be forced to peak early – as they must be high star forming at some point in order to end up with high mass in stars – and then undergo rapid quenching. Thus, it appears that a fixed present-day SM correlates to whether a galaxy exhibits quenching, as galaxies seem not to want to quench without this constraint, and a fixed present-day SFR correlates to a galaxy’s final mass bracket.
@@
Finally, we note that when comparing simulations, varying SFR(0) results in similar evolutionary tracks in the shifted analogues, but different conclusions regarding the significance of SFR as a selection parameter compared to SM alone. In EAGLE, the extreme choices in SFR(0) show drastic differences in evolutionary tracks of the shifted analogues – on the low SFR(0) end (panels (a) and (f)), the SM assembly is too quick and star formation quenches too rapidly compared to the SM-controlled group (black), and on the high SFR(0) end (panels (e) and (j)) the SM assembles too slowly and the SFR does not quench enough compared to the SM-controlled sample. However, the original MWAs (centre panels (c) and (h), green) trace closely the SM-controlled group for both the SMHs and SFHs as previously discussed, and lay in a sweet spot where present-day SFR and SM seem to correlate well, implying SFR may be a redundant selection criterion.
@@
In contrast, the extreme low end of SFR(0) (leftmost panels (k) and (p)) in TNG results in near perfect overlap between the shifted analogues (pink) and the SM-controlled group (black), while the original MWAs (centre panels (m) and (r), orange) assemble their SM too slowly and avoid strong quenching as compared to the SM-controlled group. The MWAs do not lay in the same sweet spot as in EAGLE where present-day SFR and SM seem to correlate such that MWA evolution follows that of the SM-controlled group and the inclusion of present-day SFR as a selection criterion is redundant, and as such present-day SFR is necessary to include as a parameter in MWA selection.
@@ 6.3 Re-evaluating the selection parameters
The results of our study of MWA SFHs in EAGLE and TNG produced conflicting views on the importance of present-day SFR as a MWA selection criterion. The simulations imply different conclusions in the context of MWA evolutionary tracks – in EAGLE, SM is sufficient as a stand-alone selection criterion for MWAs, but in TNG it is necessary to include SFR as a parameter. Both simulations do agree, however, that including SFR as a parameter for OSAs drives down selection efficiency.
@@
If we consider the two results that (1) according to EAGLE’s SFHs, SFR can safely be neglected as a selection parameter; and (2) according to both simulations, the SFR constraint in observational studies may not be useful, then the combination of these suggests that the inclusion of SFR as a selection parameter may be too restrictive.
@@
To test the effect of loosening the SFR constraint, we have re-calculated , as in Equation (10), using a series of expanding SFR “pseudo selection boxes”, corresponding to a change in SFR(0) uncertainty ranging from 1 to 3 times the value of 0.19 provided by LNB15 in regular intervals of 0.2. Here we remind the reader that the use of the term “pseudo” is to signify that neither the MWAs nor the OSAs are identified using a harsh selection box, but are weighted with respect to this outlined uncertainty. We also clarify that we are only expanding the range in SFR(0) used to find the true MWAs(0) (i.e., increasing the in Equation (4)) while leaving the range in SFR used to select the OSAs untouched (i.e., in Equation (8) remains the same). In other words, the “mock observer” does not change the way in which they select OSAs at ; they continue to select OSAs using the we provide them as calculated from our MWAs, and they are unaware of our loosening the SFR(0) constraint in selecting true MWAs(0).
The results are plotted in Figure 11,
@@
where increasing transparency in the line corresponds to increasing size in the SFR “pseudo selection box”. We discover that loosening the SFR(0) constraint in this way results in a greater proportion of true MWAs and a smaller proportion of MWA-“imposters” – in the sense that more OSAs satisfy the present-day SM and SFR criteria to be a MWA(0) – significantly reducing contamination at higher redshifts. Increasing the SFR(0) cutoff produces a smoother transition in going back in time, particularly for the largest selection box which produces selection nearly as efficient, and sometimes even better, than the SM-controlled group333Note that the OSA selection (translucent green/orange lines in Figure 11) will not perfectly converge to the black SM-controlled line even with an infinitely large SFR(0) “selection box” (i.e., making the SFR term in our initial MWA(0) selection vanish in Equation (4)). This is because the distribution of SFR still is present in the SFR term in the OSA selection (second term in Equation (8)). SFR-dependence only completely disappears if SFR is not considered either selection at all – neither in MWA(0) selection nor OSA selection.. Future studies of MWAs, especially in observations, may benefit from such an expanded selection window in SFR(0). We also note that an expanded selection window should have minimal consequences on our study of MWA evolution, as the effects of incorporating slightly higher and lower SFRs(0) should essentially cancel, as evident from the panels directly to the right (panels (d) and (n)) and left (panels (b) and (l)) of the MWAs’ SFHs in Figure 10.
@@
To understand why expanding the “pseudo selection box” in SFR(0) increases selection efficiency, we first explain why for the SFR-controlled group in Figure 9 is so low, which ends up being simulation-dependent. In EAGLE, instantaneous SFR tends to bounce around between snapshots due to high SFR variability in the simulation (see Figure 7), pushing the limits of the MWA distribution further out and expanding the cutoff of the OSA “pseudo selection box”. Consequently, many non-MWA(0) galaxies will get picked up as OSAs, driving down.
@@
There is a different cause for the decline in in TNG. We previously noted that for the MWAs in Figure 7 plateaus after going back Gyr, indicating that SFR variability is not as strong as in EAGLE; rather, the MWAs as a group evolve cohesively. This is illustrated in Figure 10, where all groups but the one selected on the lowest SFR(0) exhibit little dispersion in their SFHs (see panels (l) – (o)). What is causing the drop in TNG then is the overlap between these shifted analogue groups. Looking at the SFHs of the shifted analogues to either side of the MWAs, we see that these groups all occupy much of the same space. Thus, the OSA “pseudo selection box” is picking up many of these galaxies with SFR(0) slightly above or below the MW’s present-day value, hence why expanding the SFR cutoff significantly increases efficiency.
@@ 7 Summary & conclusions
Our galaxy population comes from the cosmological, hydrodynamical simulations EAGLE and TNG. We identify MWAs from this greater population using a distribution centred on the MW’s SM and SFR given by Equation (4), re-weighted against the background galaxies to offset a skew towards low SM or low SFR. This process results in a set of MW-ness weights, , given by Equation (7), providing a metric of similarity to the MW for each galaxy within the simulations. In addition to our MWAs, we use the same method to identify two control groups, one selected on SM alone and the other on SFR alone, in order to determine the role of each parameter in altering the sample of MWA evolutionary tracks.
@@
The main results of our work are summarized as follows, noting again that our results refer only to MWAs as defined by their present-day SM and SFR, and our conclusions should not be extrapolated to MWAs of any other type:
@@
MWAs’ mass assembly histories are best predicted by present-day SM. The SMHs of the MWAs in Figure 6 (top panels) follow very closely to the SM-controlled group, with near-perfect overlap between the two groups in EAGLE, indicating all galaxies with a present-day SM comparable to the MW assemble their mass in stars similarly over time, regardless of final SFR.
MWAs show a characteristic SFH profile predicted by present-day SM at early times and present-day SFR at late times, corresponding to an early peak in SFR that remains elevated through cosmic time. In Figure 6 (bottom panels), the SFHs of the MWAs align more closely with the SM-controlled group at early times, before crossing over to aligning more closely with the SFR-controlled group at late times. This crossover point, where the tracer of star formation evolution changes from SM to SFR, occurs much earlier and results in a stronger divergence in TNG than in EAGLE.
The divergence between the MWAs and SM-controlled sample mentioned above leads to another result: in TNG, MWAs do not quench as strongly as most galaxies of the same present-day SM. The SFHs in Figure 6 (bottom right panel) show that galaxies with a MW-like SM (i.e., the SM-controlled group) in TNG rapidly quench early in their evolution, but the MWAs are the ones “lucky” enough to bypass this shutoff of star formation. The reason for this stark difference when including SFR in the selection criteria for TNG, allowing them to avoid quenching, will be the focus of a future publication. We also note that this lack of quenching is why the TNG MWAs appear too blue in Figure 4, as the galaxies have always been forming new stars and hence do not turn red.
In contrast to the above point, all MW-SM galaxies in EAGLE experience slow quenching with the MWAs holding only a slightly elevated present-day SFR in comparison (see Figure 6, bottom left panel). Thus, it seems again that present-day SM is the best predictor for MWA evolution in EAGLE. In fact, we conclude that present-day SM may be sufficient as a stand-alone selection parameter for MWAs in EAGLE, however in TNG it is necessary to include present-day SFR.